x Your answer is incorrect. Try again. Columbia Sportswear Companys financial statements are presented in Appendix B. ClickCOLUMBIA SPORTSWEAR COMPANY CONSOLIDATED STATEMENTS OF OPERATIONS (In thousands, except per share amounts) Year Ended DecembeCOLUMBIA SPORTSWEAR COMPANY CONSOLIDATED STATEMENTS OF COMPREHENSIVE INCOME (In thousands) Year Ended December 31. 2016 2015COLUMBIA SPORTSWEAR COMPANY CONSOLIDATED BALANCE SHEETS (In thousands) December 31. 2016 2015 ASSETS $ 551,389 $ 369,770 4722016 2015 2014 $ 198,439 $ 179,851 $ 141.859 54,017 481 60,016 4,805 (19,178) 10.986 56,521 5,098 (11.709) 11,672 (7,873) (6,CONSOLIDATED STATEMENTS OF EQUITY (In thousands) Total S 1,252,864 141,859 10 7,751 (27,789) (39.836 19.136 5,119 11,120 (15.
Answer #1

Formula for calculating Free Cash Flow for Columbia sportswear for the year 2016

Operating cash flows - Increase in Working capital - Capital expenditures

256508- (1049172 - 883319) - 49987 = 40668

working note :

working capital = Current assets - current liabilities

Changes in Working capital = WC 2016-WC 2015

WC 2016 = 1412023 - 362851 = 1049172

WC 2015 = 1249389 - 366070 = 883319

(No other information about VF corp or appendix C is given)